OpenStudy (anonymous):

On the day that Jennifer was born, her rich uncle set up a trust for her and invested $20000 at a 8% annual interest rate compounded monthly. After how many months will the investment first be worth at least $77000? (integer answer)

OpenStudy (ahsome):

The equation you use for compound interest is: \[FV = PV × (1+r)^n\] Where: \[FV=Future Value\] \[PV=Present Value\] \[R=Rate\] \[n=time\] To find the time, use this equation \[n = ln(FV / PV) / ln(1 + r)\] Where ln=Natural logarithm. Most calculators have this button. Substitute the values \[n = ln(77000 / 20000) / ln(1 + 1.08)\] This will give you the answer, @imyint

OpenStudy (ahsome):

For more info: http://www.mathsisfun.com/money/compound-interest.html

OpenStudy (aum):

Since the interest is compounded MONTHLY (not annually), the interest rate in decimal should be divided by 12.

OpenStudy (ahsome):

^^ Yes ^^. Although this is not always the case if the rate itself is monthly

OpenStudy (aum):

Denominator: ln( 1 + 0.08/12) (not 1.08/12).

OpenStudy (ahsome):

Really? I thought it would be 1.08 as then you would get the percentage, but you probably know way more than me ;) Rewritten equation: \[n=\frac{ln(77000/20000)}{ln(1+(0.08/12))}\]

OpenStudy (aum):

Round your answer to the next highest integer because the question asks: "AT LEAST $77000? (INTEGER answer)"